OpenStudy (anonymous):

what is te equation of the line in standard form that passes through the point (-2,-3) and is parallel to the line y=x+9

OpenStudy (anonymous):

Any equation of line parallel to y=x + 9 can be given by, y= x+k-----------A Now, since the line passes through (-2, -3)... put x=-2 and y=-3 in the equation to get the value of k.. as, -3 = -2 + k or, k = -3+2 = -1 Now, put the value of k in the equation A... y=x -1.... is your required equation....

OpenStudy (anonymous):

Thanks:)